OpenStudy (anonymous):

solve: 2x^2 - x = 15

jhonyy9 (jhonyy9):

2x2 -x -15 =0 x_1,_2=(1+/- sqrt(1+120))/4 =(1+/- sqrt121)/4 = (1+/- 11)/4 = -10/4 = -5/2 and 12/4 =3

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

thanks

jhonyy9 (jhonyy9):

yw and good luck medal ?

jhonyy9 (jhonyy9):

thank you